# Loading non-Maxwellian Velocity Distributions in Particle Simulations

### v1

@@ Abstract
Numerical procedures for generating non-Maxwellian velocity distributions in particle simulations are presented. First, Monte Carlo methods for an $(r,q)$ distribution that generalizes flattop and kappa distributions are discussed. Then, two rejection methods for the regularized kappa distribution are presented, followed by a comparison in the $\kappa$ space. A simple recipe is proposed for the subtracted kappa distribution. Properties and numerical recipes for the ring and shell distributions with a finite Gaussian width are discussed. The ring and shell Maxwellians are further introduced as alternatives to the ring and shell distributions. Finally, the super-Gaussian and the filled-shell distributions are presented.
@@ I Introduction
Plasma velocity distribution functions (VDFs) show many different faces in the heliosphere. Since the mean free path is much longer than the typical length scale, local VDFs are very different from a Maxwell distribution. In the presence of a planetary dipole field, a VDF is often loss-cone-shaped, because particles with field-aligned velocities escape to the planet. The loss-cone VDF leads to pressure anisotropy, and it in turn generates various plasma waves in planetary magnetospheres. A VDF sometimes has a power-law tail in the high-energy part. Such a VDF is often approximated by a generalized Lorentzian distribution, called the kappa distribution [29, 43]. The kappa distribution has drawn significant attention in plasma physics, because it might be an outcome of long-range interaction between charged particles [21, 19, 40]. Extensions of the kappa distribution have been actively developed, such as the kappa distribution with a high-energy cutoff and the distribution with a loss-cone [35, 38, 37]. In the solar wind, electron VDFs often have kappa-type and other subcomponents [45]. In addition, ion VDFs often contain a ring- or shell-shaped “pickup” component that originates from interstellar neutral atoms [25, 57]. A shell-shaped VDF further evolves to a filled-shell VDF, due to an adiabatic expansion of the solar wind. In the presence of pickup ions (PUIs), local plasma processes can be substantially modified [23, 27, 31]. Associated with electron heating, flattop-type VDFs are often observed in and around the magnetosphere [12, 39, 32, 28, 34]. Indeed, a wide variety of VDFs are recognized throughout the heliosphere.
@@
Particle-in-cell (PIC) simulation [7, 15] is very useful to explore kinetic processes in plasma systems. Without PIC simulation, it is virtually impossible to predict wave-particle interaction, pitch-angle scattering, and particle acceleration in the aforementioned VDFs. For a Maxwell distribution, one can use a built-in procedure to generate a normal distribution or one can use the Box–Muller [8] method. However, it is not always clear how to prepare non-Maxwellian VDFs in the simulation. For example, despite its importance, numerical procedures for kappa distributions have been largely unknown to the community, until Abdul & Mace [1, 2] utilized its relation to Student’s $t$ distributions. Zenitani & Nakano [53, 54] have proposed numerical procedures for kappa, relativistic, and loss-cone distributions only recently.
@@
Yet, there are many other VDFs that simulation scientists would like to use, as will be discussed in this article. To prepare these distributions, one may consider general-purpose methods such as the acceptance-rejection sampling or the inverse-transform sampling. In the acceptance-rejection method, we have to define a good envelope function to make the rejection process efficient; otherwise the algorithm becomes increasingly inefficient, in particular in three dimensions. The inverse-transform method is based on the inverse function of the cumulative distribution function (CDF), but the inverse function is not always given analytically. Also, it is difficult to extend the method to multi dimensions. Considering limitations and uncertainties of the general-purpose methods, it is necessary to develop dedicated numerical procedures for various non-Maxwellian VDFs in heliophyiscs.
@@
The purpose of this paper is to comprehensively present numerical procedures for generating non-Maxwellian VDFs in particle simulations, that are not covered by our recent studies [53, 54]. We provide “numerical recipes” for nine different distributions, in the hope that they make kinetic modeling of plasma processes easier than before. The rest of the paper is organized as follows. In Section II, we summarize random number generators for basic probability distributions in statistics. Section III discusses the $(r,q)$ distribution that generalizes the kappa and flattop distributions. A gamma-distribution method and a piecewise rejection method are evaluated in the 2-D parameter space. Section IV discusses the regularized kappa distribution. Two rejection methods, a post-rejection method and a piecewise rejection method are proposed. Section V develops a procedure for a subtracted kappa distribution, a recently-proposed kappa loss-cone (KLC) model with a narrow loss cone. Section VI discusses ring and shell distributions with finite Gaussian width. Two methods are briefly presented. Section VII presents ring and shell Maxwellians. In particular, we introduce a shell Maxwell distribution as a handy alternative to the shell distribution. Section VIII presents other isotropic distributions. The super Gaussian distribution and the filled-shell distributions are discussed. Section IX contains discussion and summary.
II Preparation
@@ II.2 Gamma distribution
  • For $k>1$, Marsaglia & Tsang [22]’s algorithm is the de-facto standard. There are simpler methods for integer and half-integer cases (Kroese et al. [17], Yotsuji [50]. See also Zenitani & Nakano [53, Appendix A]).
@@ II.3 Generalized beta-prime distribution
A beta-prime variate can be obtained from two gamma variates [53, Section III].
@@ III $(r,q)$ distribution
The generalized $(r,q)$ distribution was proposed to model flattop-type non-Maxwellian VDFs [51, 33]. It occasionally draws attention in and around the terrestrial magnetosphere [e.g., 32, 34]. The phase-space density of the distribution is defined as follows:
@@
where $N_{0}$ is the plasma number density, $C_{\mathrm{rq}}$ is the normalization constant, and $\theta_{\parallel}$ and $\theta_{\perp}$ are characteristic speeds in the parallel and perpendicular directions with respect to the magnetic field. We consider $r\geq 0$ and $q>1$. The pressures and the kinetic energy density of the distribution are
@@
The phase-space density (Eq. (7)) with $(r,q,\theta_{\parallel},\theta_{\perp})=(2,2,\theta,\theta)$ is shown by the black line in Figure 1(a). The distribution gives a flattop-like picture.
@@
Figure 1: (a) Phase-space density of the $(r,q)$ distribution with $(r,q,\theta_{\parallel},\theta_{\perp})=(2,2,\theta,\theta)$. The distribution function (black line; Eq. (7)) and normalized Monte Carlo results (blue histogram) of $10^{6}$ particles. The solid and dashed orange lines show the left and right envelope functions (Eq. (19)). (b) Acceptance efficiency for the $(r,q)$ distribution (Eq. (22)). The vertical red line corresponds to the kappa distribution and the red dashed line corresponds to the flattop distribution. Below the blue dashed line, the beta-prime method requires a gamma distribution with a shape parameter less than unity in the denominator.
@@
Practically, one can generate the $(r,q)$ distribution by using either of the two methods. The piecewise rejection method can be easily implemented, but it becomes increasingly inefficient in a kappa-like regime in the upper-left in Figure 1(b). In such cases, the beta-prime method is favorable. Along the red line, one should use dedicated methods for the kappa distribution [2, 53]. Below the blue curve in Figure 1(b), $q\leq 1+\frac{3}{2(1+r)}$, we recommend the piecewise rejection method, because the beta-prime method may encounter the zero-division problem. Using Algorithm 3.1, we have numerically generated $10^{6}$ particles that follows the $(r,q)$ distribution with $(r,q)=(2,2)$. Their distribution is shown by the blue histogram in Figure 1(a). The numerical results are in excellent agreement with Eq. (7).
@@ IV Regularized kappa distribution
Algorithm 4.1: Post-rejection method
@@
Algorithm 4.2: Piecewise rejection method
@@
$S_{L}\leftarrow\log(1+x_{c})~~~~~~~~~~~~~~~~~~~~~~~(\mathrm{\kappa=1/2})$
@@
and then we find the following inverse-transform procedure
@@
with help from the asymptotic relation for $x_{c}\rightarrow\infty$ [30, Section 13, 13.2.18],
@@ V Subtracted kappa distribution
The kappa loss-cone (KLC) distributions have been studied for decades [38, 47]. Similarly, the loss-cone is approximated by a power of the perpendicular velocity [38] or the pitch-angle sine [47, 54], but it was very recently that Summers & Stone [37] proposed the subtracted kappa distribution, a kappa extension of the subtracted Maxwellian. This new model is promising for studying kinetic processes in the KLC distribution with a narrow loss-cone. The phase-space density $f_{\mathrm{sk}}$, its normalization constant $C_{\mathrm{sk}}$, pressures and energy density are given by:
@@
$P_{\parallel}$ $=\frac{\kappa}{2\kappa-3}N_{0}m\theta_{\parallel}^{2},~~P_{\perp}=\frac{\kappa(1+\beta[1-\Delta])}{2\kappa-3}N_{0}m\theta_{\perp}^{2},~~\mathcal{E}=\frac{\kappa}{2\kappa-3}N_{0}m\left(\frac{1}{2}\theta_{\parallel}^{2}+(1+\beta[1-\Delta])\theta_{\perp}^{2}\right)$ (51)
Here, the kappa index $\kappa>3/2$ controls a slope in the high-energy tail, the shape parameter $\beta\in[0,1]$ controls the opening angle of the loss cone, the filling parameter $\Delta$ controls the density inside the loss cone, and $\theta_{\parallel}$ and $\theta_{\perp}$ are the parallel and perpendicular characteristic velocities. When $\kappa\rightarrow\infty$, the distribution reverts to the subtracted Maxwellian. When $\beta=0$ or $\Delta=1$, it becomes a bi-kappa distribution with no loss cone. When $\beta=1$, the distribution approaches the Dory-type (Summers-type) KLC distribution with loss-cone index $j=1$ [38, $\sigma=1$ in their notation]. When $\Delta=0$, the loss cone is well developed. When $\Delta=1$, the loss cone is completely filled, and then the distribution reverts to a bi-Maxwellian.
@@
Algorithm 5: subtracted kappa
@@
The two terms correspond to the exponential distributions (Section II.1). The probability distribution of $x$ ($G(x)$) is given by their convolution [54, Section 2].
@@
Figure 3: Monte Carlo sampling of the subtracted kappa distribution ($\theta=1.0$, $\kappa=3.5$, $\beta=0.5$) with $10^{6}$ particles. Phase-space density is shown in the $_{\perp}vv_{\parallel}$ plane. The filling parameter is (a) $\Delta=0.0$ and (b) $\Delta=0.2$, respectively.
We have verified our method by a Monte Carlo test with $10^{6}$ particles. Figure 3 shows the phase-space density of the subtracted kappa distribution ($\theta=1.0$, $\kappa=3.5$, $\beta=0.5$) in the $_{\perp}vv_{\parallel}$ plane. The filling parameters are (a) $\Delta=0.0$ and (b) $\Delta=0.2$, respectively. We recognize a narrow loss-cone in the $\Delta=0.0$ case and a partially-filled loss-cone in the $\Delta=0.2$ case. Although not shown here, we have compared the numerical results and the exact solution at several 1-D cuts in the velocity space, and found that they are in excellent agreement. We have also conducted a Kullback–Leibler (KL) divergence test [54], to check the numerical result.
@@ VI Ring and Shell distributions
We first discuss a ring distribution with a Gaussian width [e.g., 46, 18]. We show the phase-space density $f_{\mathrm{ring}}(\bm{v})$, its perpendicular part $g_{\mathrm{ring}}(v_{\perp})$, and an auxiliary function $A_{2}(x)$.
$f_{\mathrm{ring}}(\bm{v})d^{3}{v}$ $=N_{0}\frac{1}{{\pi}^{3/2}\theta_{\parallel}\theta_{\perp}^{2}A_{2}(\frac{V}{\theta_{\perp}})}\exp\left(-\frac{v_{\parallel}^{2}}{\theta_{\parallel}^{2}}-\frac{(v_{\perp}-V)^{2}}{\theta_{\perp}^{2}}\right)d^{3}{v},$ (67)
@@
$g_{\mathrm{ring}}({v}_{\perp})$ $=\frac{2v_{\perp}}{\theta_{\perp}^{2}A_{2}(\frac{V}{\theta_{\perp}})}\exp\left(-\frac{(v_{\perp}-V)^{2}}{\theta_{\perp}^{2}}\right),$ (69)
$A_{2}(x)$ $=\exp(-x^{2})+\sqrt{\pi}x\,\mathrm{erfc}\left(-x\right),$ (70)
Here, $V$ is the ring velocity, $\theta_{\parallel}$ and $\theta_{\perp}$ are parallel and perpendicular characteristic velocities, and $\mathrm{erfc}(x)$ is the complementary error function. The perpendicular and parallel components are independent. The parallel component is assumed to be a Maxwellian. In the context of the solar wind, $V$ corresponds to the solar-wind speed in the perpendicular direction. The pressures and the energy density are given by [18]
$P_{\perp}$ $=\frac{N_{0}m\theta_{\perp}^{2}}{4A_{2}(\frac{V}{\theta_{\perp}})}\left\{2\left(1+\frac{V^{2}}{\theta_{\perp}^{2}}\right)\exp\left(-\frac{V^{2}}{\theta_{\perp}^{2}}\right)+\sqrt{\pi}\frac{V}{\theta_{\perp}}\left(3+2\frac{V^{2}}{\theta_{\perp}^{2}}\right)\,\mathrm{erfc}\left(-\frac{V}{\theta_{\perp}}\right)\right\}$ (71)
$=\frac{N_{0}m\theta_{\perp}^{2}}{2}\left\{1+\frac{V^{2}}{\theta_{\perp}^{2}}+\frac{\sqrt{\pi}}{2A_{2}(\frac{V}{\theta_{\perp}})}\frac{V}{\theta_{\perp}}\mathrm{erfc}\left(-\frac{V}{\theta_{\perp}}\right)\right\}$ (72)
$P_{\parallel}$ $=\frac{1}{2}N_{0}m\theta_{\parallel}^{2},~~~~\mathcal{E}=\frac{1}{4}N_{0}m\left\{\theta_{\parallel}^{2}+{2\theta_{\perp}^{2}}\left[1+\frac{V^{2}}{\theta_{\perp}^{2}}+\frac{\sqrt{\pi}}{2A_{2}(\frac{V}{\theta_{\perp}})}\frac{V}{\theta_{\perp}}\mathrm{erfc}\left(-\frac{V}{\theta_{\perp}}\right)\right]\right\}$ (73)
When $V=0$, the distribution reverts to a bi-Maxwellian.
We also discuss the shell distribution with a Gaussian width [e.g., 13, 49, 20, 24]. We show the phase-space density $f_{\mathrm{shell}}(\bm{v})$, its radial form $g_{\mathrm{shell}}(v)$, and an auxiliary function $A_{3}(x)$.
$f_{\mathrm{shell}}(\bm{v})d^{3}{v}$ $=\frac{N_{0}}{2\pi\theta^{3}A_{3}(\frac{V}{\theta})}\exp\left(-\frac{(|v|-V)^{2}}{\theta^{2}}\right)d^{3}{v}=N_{0}\bigg\{g_{\mathrm{shell}}({v})dv\bigg\},$ (74)
$g_{\mathrm{shell}}({v})$ $=\frac{2v^{2}}{\theta^{3}A_{3}(\frac{V}{\theta})}\exp\left(-\frac{(|v|-V)^{2}}{\theta^{2}}\right),$ (75)
$A_{3}(x)$ $=x\exp({-x^{2}})+\sqrt{\pi}\left(x^{2}+\frac{1}{2}\right)\mathrm{erfc}{\left(-x\right)}.$ (76)
@@
${P}$ $=\frac{N_{0}m\theta^{2}}{12A_{3}(\frac{V}{\theta})}\left\{2\left(5+2\frac{V^{2}}{\theta^{2}}\right)\frac{V}{\theta}\exp\left(-\frac{V^{2}}{\theta^{2}}\right)+\sqrt{\pi}\left(3+12\frac{V^{2}}{\theta^{2}}+4\frac{V^{4}}{\theta^{4}}\right)\mathrm{erfc}\left(-\frac{V}{\theta}\right)\right\}$ (77)
$=\frac{N_{0}m\theta^{2}}{6}\left\{5+2\frac{V^{2}}{\theta^{2}}-\frac{\sqrt{\pi}}{A_{3}(\frac{V}{\theta})}\mathrm{erfc}\left(-\frac{V}{\theta}\right)\right\}$ (78)
$\mathcal{E}$ $=\frac{N_{0}m\theta^{2}}{4}\left\{5+2\frac{V^{2}}{\theta^{2}}-\frac{\sqrt{\pi}}{A_{3}(\frac{V}{\theta})}\mathrm{erfc}\left(-\frac{V}{\theta}\right)\right\}$ (79)
@@
$\mathrm{CDF}_{\mathrm{ring}}(v_{\perp})$ $=1-\frac{\exp\left(-\frac{(v_{\perp}-V)^{2}}{\theta_{\perp}^{2}}\right)+\sqrt{\pi}\frac{V}{\theta_{\perp}}\,\mathrm{erfc}\left(\frac{v_{\perp}-V}{\theta_{\perp}}\right)}{A_{2}(\frac{V}{\theta})}$ (80)

### v2

@@ Abstract
Numerical procedures for generating non-Maxwellian velocity distributions in particle simulations are presented. First, Monte Carlo methods for an $(r,q)$ distribution that generalizes flattop and Kappa distributions are discussed. Then, two rejection methods for the regularized Kappa distribution are presented, followed by a comparison in the $\kappa$ space. A simple recipe is proposed for the subtracted Kappa distribution. Properties and numerical recipes for the ring and shell distributions with a finite Gaussian width are discussed. The ring and shell Maxwellians are further introduced as alternatives to the ring and shell distributions. Finally, methods for the super-Gaussian and the filled-shell distributions are presented.
@@ I Introduction
Plasma velocity distribution functions (VDFs) show many different shapes in the heliosphere. Since the mean free path is much longer than the typical length scale, local VDFs deviate significantly from a Maxwell distribution. In the presence of a planetary dipole field, the VDF is often loss-cone-shaped, because particles with field-aligned velocities escape to the planet. The loss-cone VDF leads to pressure anisotropy, which in turn generates various plasma waves in planetary magnetospheres. VDFs sometimes exhibit a power-law tail in the high-energy part. Such VDFs are often approximated by a generalized Lorentzian distribution, now widely known as the Kappa distribution [31, 46]. The Kappa distribution has drawn significant attention in plasma physics, because it might be an outcome of long-range interaction between charged particles [22, 20, 43]. Extensions of the Kappa distribution have been proposed and actively studied, such as the Kappa distribution with a high-energy cutoff and the distribution with a loss-cone [38, 41, 40]. In the solar wind, electron VDFs often have Kappa-type and other subcomponents [48]. In addition, ion VDFs often contain a ring- or shell-shaped “pickup” component that originates from interstellar neutral atoms [27, 62]. A shell-shaped VDF further evolves to a filled-shell VDF, due to an adiabatic expansion of the solar wind. In the presence of pickup ions (PUIs), local plasma processes can be substantially modified [24, 29, 33]. Associated with electron heating, flattop-type VDFs are often observed in and around the magnetosphere [12, 42, 34, 30, 37]. Indeed, a wide variety of VDFs are recognized throughout the heliosphere.
@@
Particle-in-cell (PIC) simulation [7, 15] is very useful for exploring kinetic processes in plasma systems. Without PIC simulation, it is virtually impossible to accurately predict wave-particle interaction, pitch-angle scattering, and particle acceleration in the aforementioned VDFs. For a Maxwell distribution, one can use a built-in procedure to generate a normal distribution or one can use the Box–Muller [8] method. However, it is not always clear how to generate non-Maxwellian VDFs in simulation. For example, despite its importance, numerical procedures for Kappa distributions have been largely unknown to the community, until Abdul & Mace [1, 2] used its relation to Student’s $t$ distributions. More recently, Zenitani & Nakano [56, 57] proposed numerical procedures for Kappa, relativistic, and loss-cone distributions. Their methods rely on only three types of basic random numbers: uniform, normal, and gamma random numbers.
@@
Yet, there are many other VDFs that simulation scientists would like to use, as discussed later in this article. To generate such distributions, one may consider general-purpose methods such as acceptance-rejection sampling or inverse transform sampling. In the acceptance-rejection method, we need to define a good envelope function to make the rejection process efficient; otherwise the algorithm becomes increasingly inefficient, especially in three dimensions. The inverse transform method relies on the inverse of the cumulative distribution function (CDF), which is not always available analytically. Moreover, it is difficult to extend the method to multiple dimensions. Considering the limitations of these general-purpose methods, it is necessary to develop dedicated numerical procedures for a variety of non-Maxwellian VDFs in heliophysics.
@@
The purpose of this paper is to comprehensively present numerical procedures for generating non-Maxwellian VDFs in particle simulations, that are not covered by our recent studies [56, 57]. We provide “numerical recipes” for nine different distributions, in the hope that they make kinetic modeling of plasma processes easier than before. The rest of the paper is organized as follows. In Section II, we summarize random number generators for basic probability distributions in statistics. Section III discusses the $(r,q)$ distribution that generalizes the Kappa and flattop distributions. A gamma-distribution method and a piecewise rejection method are evaluated in the 2-D parameter space. Section IV discusses the regularized Kappa distribution. Two rejection methods, a post-rejection method and a piecewise rejection method are proposed. Section V develops a procedure for a subtracted Kappa distribution, a recently-proposed Kappa loss-cone (KLC) model with a narrow loss cone. Section VI discusses ring and shell distributions with finite Gaussian width. Two methods are briefly presented. Section VII presents ring and shell Maxwellians. In particular, we introduce a shell Maxwell distribution as a handy alternative to the conventional shell distribution. Section VIII presents other isotropic distributions. The super Gaussian distribution and the filled-shell distributions are discussed. Section IX validates numerical accuracies of the proposed methods for all the distributions. Section X contains a discussion and summary.
II Basic statistical distributions
@@ II.2 Gamma distribution
  • For $k>1$, Marsaglia & Tsang [23]’s algorithm is the de-facto standard. There are simpler methods for integer and half-integer cases (Kroese et al. [18], Yotsuji [53]. See also Zenitani & Nakano [56, Appendix A]).
@@ II.3 Generalized beta-prime distribution
A beta-prime variate can be obtained from two gamma variates [56, Section III].
@@ III $(r,q)$ distribution
The generalized $(r,q)$ distribution was proposed to model flattop-type non-Maxwellian VDFs [54, 36]. It often draws attention in and around the terrestrial magnetosphere [e.g., 34, 35, 37]. The phase space density of the distribution is defined as follows:
@@
where $N_{0}$ is the plasma number density, $C_{\mathrm{rq}}$ is the normalization constant, and $\theta_{\parallel}$ and $\theta_{\perp}$ are characteristic speeds in the parallel and perpendicular directions with respect to the magnetic field. We consider r>1r>-1 and $q>1$. The pressures and the kinetic energy density of the distribution are
@@
The phase space density (Eq. (7)) with $(r,q,\theta_{\parallel},\theta_{\perp})=(2,2,\theta,\theta)$ is shown by the black line in Figure 1(a). The distribution gives a flattop-like picture.
@@
Figure 1: (a) Phase space density of the $(r,q)$ distribution with $(r,q,\theta_{\parallel},\theta_{\perp})=(2,2,\theta,\theta)$. The distribution function (black line; Eq. (7)) and normalized Monte Carlo results (blue histogram) of $10^{6}$ particles. The solid and dashed orange lines show the left and right envelope functions (Eq. (19)). (b) Acceptance efficiency for the $(r,q)$ distribution (Eq. (22)). The vertical red line corresponds to the Kappa distribution and the red dashed line corresponds to the flattop distribution. Below the blue dashed line, the beta-prime method requires a gamma distribution with a shape parameter less than unity in the denominator.
@@
Practically, one can generate the $(r,q)$ distribution by using either of the two methods. The piecewise rejection method can be easily implemented, but it becomes increasingly inefficient in a Kappa-like regime in the upper-left in Figure 1(b). In such cases, the beta-prime method is favorable. Along the red line, one should use dedicated methods for the Kappa distribution [2, 56, 60]. Below the blue curve in Figure 1(b), $q\leq 1+\frac{3}{2(1+r)}$, we recommend the piecewise rejection method, because the beta-prime method may encounter the zero-division problem. Using Algorithm 3.1, we have numerically generated $10^{6}$ particles that follows the $(r,q)$ distribution with $(r,q)=(2,2)$. Their distribution is shown by the blue histogram in Figure 1(a). The numerical results are in excellent agreement with Eq. (7).
@@ IV Regularized Kappa distribution
Algorithm 4.1: Post-rejection method ($\kappa>1/2$)
@@
Algorithm 4.2: Piecewise rejection method ($\kappa>0$)
@@
$S_{L}\leftarrow\log(1+x_{c})~~~~~~~~~~~~~~~~~~~~~~~~(\mathrm{\kappa=1/2})$
@@
and then we find the following inverse transform procedure
@@
with help from the asymptotic relation for $x_{c}\rightarrow\infty$ [32, Section 13, 13.2.18],
@@ V Subtracted Kappa distribution
The Kappa loss-cone (KLC) distributions have been studied for decades [41, 50]. Similarly, the loss-cone is approximated by a power of the perpendicular velocity [41] or the pitch-angle sine [50, 57], but it was very recently that Summers & Stone [40] proposed the subtracted Kappa distribution, a Kappa extension of the subtracted Maxwellian. This new model is promising for studying kinetic processes in the KLC distribution with a narrow loss-cone. The phase space density $f_{\mathrm{sk}}$, its normalization constant $C_{\mathrm{sk}}$, pressures, energy density, and the pressure anisotroy $A$ are given by:
@@
$P_{\parallel}$ $=\frac{\kappa}{2\kappa-3}N_{0}m\theta_{\parallel}^{2},~~P_{\perp}=\frac{\kappa(1+\beta[1-\Delta])}{2\kappa-3}N_{0}m\theta_{\perp}^{2},$ (51)
$\mathcal{E}$ $=\frac{\kappa}{2\kappa-3}N_{0}m\left(\frac{1}{2}\theta_{\parallel}^{2}+(1+\beta[1-\Delta])\theta_{\perp}^{2}\right),$ (52)
$A$ $\equiv\frac{P_{\perp}}{P_{\parallel}}-1=(1+\beta[1-\Delta])\frac{\theta_{\perp}^{2}}{\theta_{\parallel}^{2}}-1$ (53)
Here, the kappa index $\kappa>3/2$ controls a slope in the high-energy tail, the shape parameter $\beta\in[0,1]$ controls the opening angle of the loss cone, the filling parameter $\Delta$ controls the density inside the loss cone, and $\theta_{\parallel}$ and $\theta_{\perp}$ are the parallel and perpendicular characteristic velocities. When $\kappa\rightarrow\infty$, the distribution reverts to the subtracted Maxwellian. When $\beta=0$ or $\Delta=1$, it becomes a bi-Kappa distribution with no loss cone. When $\beta=1$, the distribution approaches the Dory-type (Summers-type) KLC distribution with loss-cone index $j=1$ [41, $\sigma=1$ in their notation]. When $\Delta=0$, the loss cone is well developed. When $\Delta=1$, the loss cone is completely filled, and then the distribution reverts to a bi-Maxwellian.
@@
Algorithm 5: subtracted Kappa ($\kappa>3/2$)
@@
The two terms correspond to the exponential distributions (Section II.1). The probability distribution of $x$ ($G(x)$) is given by their convolution [57, Section 2].
@@
Figure 3: Monte Carlo sampling of the subtracted Kappa distribution ($\theta=1.0$, $\kappa=3.5$, $\beta=0.5$) with $10^{6}$ particles. Phase space density is shown in the $_{\perp}vv_{\parallel}$ plane. The filling parameter is (a) $\Delta=0.0$ and (b) $\Delta=0.2$, respectively.
We have verified our method by a Monte Carlo test with $10^{6}$ particles. Figure 3 shows the phase space density of the subtracted Kappa distribution ($\theta=1.0$, $\kappa=3.5$, $\beta=0.5$) in the $_{\perp}vv_{\parallel}$ plane. The filling parameters are (a) $\Delta=0.0$ and (b) $\Delta=0.2$, respectively. We recognize a narrow loss-cone in the $\Delta=0.0$ case and a partially-filled loss-cone in the $\Delta=0.2$ case. Although not shown here, we have compared the numerical results and the exact solution at several 1-D cuts in the velocity space, and found that they are in excellent agreement.
@@ VI Ring and Shell distributions
We first discuss a ring distribution with a Gaussian width [e.g., 49, 19]. We show the phase space density $f_{\mathrm{ring}}(\bm{v})$, its perpendicular part $g_{\mathrm{ring}}(v_{\perp})$, and an auxiliary function $C_{2}(x)$.
$f_{\mathrm{ring}}(\bm{v})d^{3}{v}$ $=N_{0}\frac{1}{{\pi}^{3/2}\theta_{\parallel}\theta_{\perp}^{2}C_{2}(\frac{V}{\theta_{\perp}})}\exp\left(-\frac{v_{\parallel}^{2}}{\theta_{\parallel}^{2}}-\frac{(v_{\perp}-V)^{2}}{\theta_{\perp}^{2}}\right)d^{3}{v},$ (69)
@@
$g_{\mathrm{ring}}({v}_{\perp})$ $=\frac{2v_{\perp}}{\theta_{\perp}^{2}C_{2}(\frac{V}{\theta_{\perp}})}\exp\left(-\frac{(v_{\perp}-V)^{2}}{\theta_{\perp}^{2}}\right),$ (71)
$C_{2}(x)$ $=\exp(-x^{2})+\sqrt{\pi}x\,\mathrm{erfc}\left(-x\right),$ (72)
Here, $V$ is the ring velocity, $\theta_{\parallel}$ and $\theta_{\perp}$ are parallel and perpendicular characteristic velocities, and $\mathrm{erfc}(x)$ is the complementary error function. The perpendicular and parallel components are independent. The parallel component is assumed to be a Maxwellian. In the context of the solar wind, $V$ corresponds to the solar-wind speed in the perpendicular direction. The pressures, the energy density, and the pressure anisotropy are given by [19]
$P_{\perp}$ $=\frac{N_{0}m\theta_{\perp}^{2}}{4C_{2}(\frac{V}{\theta_{\perp}})}\left\{2\left(1+\frac{V^{2}}{\theta_{\perp}^{2}}\right)\exp\left(-\frac{V^{2}}{\theta_{\perp}^{2}}\right)+\sqrt{\pi}\frac{V}{\theta_{\perp}}\left(3+2\frac{V^{2}}{\theta_{\perp}^{2}}\right)\,\mathrm{erfc}\left(-\frac{V}{\theta_{\perp}}\right)\right\}$ (73)
$=\frac{N_{0}m\theta_{\perp}^{2}}{2}\left\{1+\frac{V^{2}}{\theta_{\perp}^{2}}+\frac{\sqrt{\pi}}{2C_{2}(\frac{V}{\theta_{\perp}})}\frac{V}{\theta_{\perp}}\mathrm{erfc}\left(-\frac{V}{\theta_{\perp}}\right)\right\}$ (74)
$P_{\parallel}$ $=\frac{1}{2}N_{0}m\theta_{\parallel}^{2},~~~~\mathcal{E}=\frac{1}{4}N_{0}m\left\{\theta_{\parallel}^{2}+{2\theta_{\perp}^{2}}\left[1+\frac{V^{2}}{\theta_{\perp}^{2}}+\frac{\sqrt{\pi}}{2C_{2}(\frac{V}{\theta_{\perp}})}\frac{V}{\theta_{\perp}}\mathrm{erfc}\left(-\frac{V}{\theta_{\perp}}\right)\right]\right\}$ (75)
$A$ $=\frac{\theta_{\perp}^{2}}{\theta_{\parallel}^{2}}\left\{1+\frac{V^{2}}{\theta_{\perp}^{2}}+\frac{\sqrt{\pi}}{2C_{2}(\frac{V}{\theta_{\perp}})}\frac{V}{\theta_{\perp}}\mathrm{erfc}\left(-\frac{V}{\theta_{\perp}}\right)\right\}-1.$ (76)
We also discuss the shell distribution with a Gaussian width [e.g., 13, 52, 21, 25]. We show the phase space density $f_{\mathrm{shell}}(\bm{v})$, its radial form $g_{\mathrm{shell}}(v)$, and an auxiliary function $C_{3}(x)$.
$f_{\mathrm{shell}}(\bm{v})d^{3}{v}$ $=\frac{N_{0}}{2\pi\theta^{3}C_{3}(\frac{V}{\theta})}\exp\left(-\frac{(|v|-V)^{2}}{\theta^{2}}\right)d^{3}{v}=N_{0}\bigg\{g_{\mathrm{shell}}({v})dv\bigg\},$ (77)
$g_{\mathrm{shell}}({v})$ $=\frac{2v^{2}}{\theta^{3}C_{3}(\frac{V}{\theta})}\exp\left(-\frac{(|v|-V)^{2}}{\theta^{2}}\right),$ (78)
$C_{3}(x)$ $=x\exp({-x^{2}})+\sqrt{\pi}\left(x^{2}+\frac{1}{2}\right)\mathrm{erfc}{\left(-x\right)}.$ (79)
@@
${P}$ $=\frac{N_{0}m\theta^{2}}{12C_{3}(\frac{V}{\theta})}\left\{2\left(5+2\frac{V^{2}}{\theta^{2}}\right)\frac{V}{\theta}\exp\left(-\frac{V^{2}}{\theta^{2}}\right)+\sqrt{\pi}\left(3+12\frac{V^{2}}{\theta^{2}}+4\frac{V^{4}}{\theta^{4}}\right)\mathrm{erfc}\left(-\frac{V}{\theta}\right)\right\}$ (80)
$=\frac{N_{0}m\theta^{2}}{6}\left\{5+2\frac{V^{2}}{\theta^{2}}-\frac{\sqrt{\pi}}{C_{3}(\frac{V}{\theta})}\mathrm{erfc}\left(-\frac{V}{\theta}\right)\right\}$ (81)
$\mathcal{E}$ $=\frac{N_{0}m\theta^{2}}{4}\left\{5+2\frac{V^{2}}{\theta^{2}}-\frac{\sqrt{\pi}}{C_{3}(\frac{V}{\theta})}\mathrm{erfc}\left(-\frac{V}{\theta}\right)\right\}$ (82)
@@
$\mathrm{CDF}_{\mathrm{ring}}(v_{\perp})$ $=1-\frac{\exp\left(-\frac{(v_{\perp}-V)^{2}}{\theta_{\perp}^{2}}\right)+\sqrt{\pi}\frac{V}{\theta_{\perp}}\,\mathrm{erfc}\left(\frac{v_{\perp}-V}{\theta_{\perp}}\right)}{C_{2}(\frac{V}{\theta})}$ (83)